OpenStudy (anonymous):

Ram, Shiv and Ganesh assemble for a contributory party. Ram brings 3 apples while Shiv brings 5. Since Ganesh did not have any apples, he contributed Rs. 8. How many rupees should Ram and Shiv respectively get, assuming each of the three consumes an equal portion of the apples? 1. 1, 7 2. 3, 5 3. 5, 3 4. 2, 6

OpenStudy (anonymous):

I'm not absolutely sure. but I would go for D.

OpenStudy (anonymous):

2,6 u mean ? how ?

OpenStudy (anonymous):

Well there's not enough for the 3 of them to get the same amount (assuming the apples are the same size) so the closest one is D.

OpenStudy (anonymous):

But like I said, I'm not sure.

OpenStudy (anonymous):

@iGreen. can you help a minute?

OpenStudy (anonymous):

umm it can be 1,7 also then

OpenStudy (anonymous):

Yes, it could. Would you only want one apple if you knew there were more?

OpenStudy (anonymous):

It's money not apple

OpenStudy (anonymous):

That is what confused me xD

OpenStudy (anonymous):

Well then I have no idea XD :P

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

can u get sm1 to help me out

OpenStudy (anonymous):

so it matches with which option ?

OpenStudy (mathmath333):

hmm question is complex

OpenStudy (perl):

I think you have to consider that they should be paid for what two people actually contribute

OpenStudy (mathmath333):

\(\large \begin{align} \color{black}{\normalsize \text{assuming each of the three consumes an equal portion of the apple }\hspace{.33em}\\~\ \normalsize \text{so each would consume }\dfrac{8}{3} \normalsize \text{part of apple }\hspace{.33em}\\~\\ \normalsize \text{this also means that ganesh consumed } \dfrac{8}{3} \normalsize \text{part of apple }\hspace{.33em}\\~\\ \normalsize \text{he paid 8 rs for } \dfrac{8}{3}~~ apples\hspace{.33em}\\~\\ \normalsize \text{price of one apple }= \dfrac{1}{3}~~rs \hspace{.33em}\\~\\ \normalsize \text{ram brought 3 apples, price of 3 apples }=\dfrac{1}{3}\times 3=1~~rs \hspace{.33em}\\~\\ \normalsize \text{shiv brought 5 apples, price of 5 apples } \dfrac{1}{3}\times 5=\dfrac{5}{3}~~rs \hspace{.33em}\\~\\ \normalsize \text{ganesh brought 0 apples, but he paid } =8~~rs \hspace{.33em}\\~\\ }\end{align}\)

OpenStudy (perl):

After dividing the 8 apples among the 3 people, each person eats 8/3 apples, but Ganesh wants to give the money to the apple owners in the proportion to which they actually shared with others. Now the first person actually contributes to other people 3 - 8/3 apples. The second person actually contributes 5 - 8/3 apples. 3 - 8/3 = 1/3 5 - 8/3 = 7/3 the ratio 1/3 : 7/3 is equivalent to 1:7 So they should be given the rupees in the ratio 1 : 7

OpenStudy (mathmath333):

@aum

OpenStudy (mathmath333):

corrected \(\large \begin{align} \color{black}{\normalsize \text{assuming each of the three consumes an equal portion of the apple }\hspace{.33em}\\~\ \normalsize \text{so each would consume }\dfrac{8}{3} \normalsize \text{part of apple }\hspace{.33em}\\~\\ \normalsize \text{this also means that ganesh consumed } \dfrac{8}{3} \normalsize \text{part of apple }\hspace{.33em}\\~\\ \normalsize \text{he paid 8 rs for } \dfrac{8}{3}~~ apples\hspace{.33em}\\~\\ \normalsize \text{price of one apple }= 3~~rs \hspace{.33em}\\~\\ \normalsize \text{ram brought 3 apples, price of 3 apples }=3\times 3=9~~rs \hspace{.33em}\\~\\ \normalsize \text{shiv brought 5 apples, price of 5 apples } 3\times 5=15~~rs \hspace{.33em}\\~\\ \normalsize \text{ganesh brought 0 apples, but he paid } =8~~rs \hspace{.33em}\\~\\ \normalsize \text{we have two equations } 1.)~~r+s=8\hspace{.33em}\\~\\ 2.)~~\dfrac{r}{s}=\dfrac{9}{15}=\dfrac{3}{5}\hspace{.33em}\\~\\ \dfrac{r+s}{s}=\dfrac{3+5}{5}\hspace{.33em}\\~\\ \dfrac{8}{s}=\dfrac{8}{5}\hspace{.33em}\\~\\ s=5 \hspace{.33em}\\~\\ r=3 \hspace{.33em}\\~\\ }\end{align}\)

OpenStudy (mathmath333):

i hope this is correct

OpenStudy (mathmath333):

still unsure